### v1

# A Typology of Decision-Making Tasks for Visualization

Abstract
Despite decision-making being a vital goal of data visualization, little work has been done to differentiate the decision-making tasks within our field. While visualization task taxonomies and typologies exist, they are often too granular for describing complex decision goals and decision-making processes, thus limiting their potential use in designing decision-support tools. In this paper, we contribute a typology of decision-making tasks that were iteratively refined from a list of design goals distilled from a literature review. Our typology is concise and consists of only three tasks: choose, activate, and create. Originally proposed by the scientific community, we extend and provide definitions for these tasks that are suitable for the visualization community. Our proposed typology offers two benefits. First, it facilitates the composition of decisions using these three tasks, allowing for flexible and clear descriptions across varying complexities and domains. Second, diagrams created using this typology encourage productive discourse between visualization designers and domain experts by abstracting the intricacies of data, thereby promoting clarity and rigorous analysis of decision-making processes. We motivate the use of our typology through four case studies and demonstrate the benefits of our approach through semi-structured interviews conducted with experienced members of the visualization community, comprising academic and industry experts, who have contributed to developing or publishing decision support systems for domain experts. Our interviewees composed diagrams using our typology to delineate the decision-making processes that drive their decision-support tools, demonstrating its descriptive capacity and effectiveness.
keywords:
decision-making, typology
[…]
This figure shows the three decision tasks in our typology: CHOOSE, ACTIVATE, and CREATE. The CHOOSE task assesses options and outputs a subset deemed optimal or best. The ACTIVATE task represents a decision where options are evaluated, and only those that meet or exceed a threshold are returned. The CREATE task represents decisions on assembling, synthesizing, or generating new information. Together, these three decision tasks can be composed to represent complex decision-making problems.
1 Introduction
Visualization researchers and practitioners alike have suggested that one of the primary goals of visualization systems is to help users make decisions[55, 61, 67, 23]. Many works have come out of the visualization community that have succeeded in helping users make important decisions (e.g., [1, 70, 16, 2, 26, 29], to name a few). However, when discussing these works at the abstract level, the taxonomies within the visualization community (e.g., [3, 74, 13, 23, 24, 48]) do not adequately provide for the complexities embedded within decision-making. Thus, they limit potential use for (1) analyzing and reflecting upon design, (2) helping to facilitate communication and design, and (3) categorizing and organizing publications on decision-making systems.
In this paper, we build upon this past work, most notably Dimara et al.[23, 24, 48, 47], and present a typology of decision-making for visualization. It consists of three decision tasks that represent specific decision problems: CHOOSE, ACTIVATE, and CREATE. The CHOOSE task assesses alternatives and outputs a subset deemed optimal or best, the ACTIVATE task represents a decision where only the inputs that meet or exceed a threshold are returned, and the CREATE task represents decisions that assemble, synthesize, or generate new information. The most critical property of this typology is that it allows us to represent the complex structures of real-world decision-making problems, as our tasks can be composed or decomposed into other tasks.
To inform our typology, we gathered design goals from a thorough literature review of decision-support tools within the visualization community. We built upon previously curated corpora[48, 57], focusing our search on papers with domain expert evaluations. We then followed an inductive coding process to identify each paper’s primary decision-making goals and processes. We then took these codes and distilled the design goals for our decision-making typology, including iterativeness, composability, generalizability, simplicity, and others. From the decision-making theory[28, 58] and our design goals, we worked towards an initial set of tasks. We then used our survey to hone and iterate on our tasks until they could accurately characterize the works included. To illustrate the abovementioned properties, we present four case studies using our typology to describe users’ decisions using decision-support tools.
We validated our typology by conducting semi-structured interviews. Our nine participants were visualization experts from academia, government laboratories, and industry who had authored one or more decision-support tools. We demonstrated our typology to the participants and asked them to apply it to their systems and answer open-ended questions about the typology, its properties, and how it compares to other typologies.
Overall, the result of our validation can be divided into three broad categories: the correctness of the typology, the utility of the typology, and how the typology compared to existing task taxonomies and typologies. The correctness of the typology can be further broken down into the consideration of completeness and expressiveness. In our interview study, participants found that the typology is complete regarding the three decision tasks and that they could successfully express their decision-making problems in diagrams using the typology. Additionally, participants found the typology to be helpful when analyzing and reflecting on the design process of visualization tools, believed that it could be useful when communicating with folks outside of visualization and that it would be able to categorize visualization decision-making publications. Finally, participants compared the typology to existing work and were excited by our typology’s level of abstraction and flexibility.
In summary, the major contributions of this paper are:
  • Design goals for decision-making typologies derived from the thematic analysis of a comprehensive literature review.
  • A typology of decision tasks, which provides an abstract yet precise way to discuss decision-making activities in visualization.
  • Four case studies in which our typology is used to describe users’ decisions in the context of decision-support tools that demonstrate the properties of our typology.
  • A semi-structured interview study that serves as an evaluation of our typology and provides a roadmap for future work.
2 Background & Related Work
Decision making is a broad topic studied across many disciplines with visualization researchers emphasizing that supporting decision making is a vital goal of data visualization [23, 35, 4, 15, 5]. The visualization community tends to focus on decision making from three broad lenses: cognition and perception, decision process support, and domain-specific decision support systems. We review the relevant literature from these areas and discuss the significance to this work.
2.1 Decision Making & Visualization
In a cognitive context, decision-making is often framed in terms of rationality and judgement[51]. In this view, the most rational decisions are ones that maximize expectation of a function[27]. However, a number of human decisions are not rational[60], a fact that has impacted the visualization community in two significant ways. The first is an increase in computational decision-making support[64, 35, 73]. While promising, it is unclear and difficult to determine under what conditions or kinds of decisions users will accept help from or implement decisions of computational models or agents[49, 29] The second impact is the study of cognitive bias and perception in visualization contexts. Cognitive biases explain irrational decision-making as the result of using heuristic functions to evaluate information in a decision-making cognitive process[22].
The occurrence of specific kinds of judgment and decision-making errors can then be traced to specific heuristic functions, and recent studies have investigated the overlap of specific biases with visualization design[20]. There is evidence that visualization design can impact the strength of a cognitive bias[21], and there exists a significant body of work that assesses the role of visual perception in judgment and decision-making[34, 51, 47]. For example, in Padilla et al.[50], the authors find a significant relationship between participants’ risk assessment and uncertainty visualization design choices. While important, cognitive and perceptual approaches focus on reasoning and judgement, rather than describing decision-making problems in domain agnostic contexts for meaningful discussions on visual support.
A second body of work in the visualization community is concerned with designing visualizations that support specific problem-solving structures for a decision-making problem. These visualization systems are unique in that they support a specific approach for decision-making rather than focusing on a specific application or domain[52, 30, 63, 6, 48]. One such decision process addressed in the community is multiple criteria decision making (MCDM) which is a mathematical formalism that enables selection of a solution given multiple competing criteria and a decision maker’s preferences for various attributes of a solution[12]. MCDM lends itself nicely to computation and is particularly useful for ranking solutions[32], defining (and therefore exploring) trade-offs among solutions[52], and the explicit use of a person’s preferences can be an invaluable source of information for recommendation systems or group decision-making[33]. However, a prescribed structure to a decision-making problem is not always appropriate, and transforming a decision-making problem to fit into a particular structure can itself be challenging for users. For example, MCDM hinges on the specification of criteria and evaluation functions for those criteria[12]. However, for some problems the selection of criteria and evaluation function is in itself difficult[52, 39, 46], and externalizing preferences as abstract values can be a hard interpretation problem for users[41]. Focusing solely on describing and discussing problems from a specific structure unnecessarily constrains possible visualization solutions, and does not capture the diverse ways people can approach complex decision-making problems.
Finally, there exists a large body of work in the visual analytics community which addresses decision-making from domain-specific contexts. For example, many decision support systems exist to address medical decisions[40, 62, 2], security and fraud detection[65, 16, 45, 56], city planning[7, 75, 26], and other domain-specific applications[54, 8, 59, 69, 11]. When supporting decision making within a particular domain, it is indeed important to take into account the nuances relevant to that domain. However, domain-specific approaches do not necessarily generalize, and do not allow for a discussion of decision-making problems across domains.
2.2 Visualization Typologies
Figure 1: Our processes for: (A) generating a list of design goals from a corpus and (B) using the design goals to create our typology.
Decision-making happens at many different levels of interaction with data and within visualization systems. Yet, no existing visualization typology can capture both high-level decisions (e.g., what car to buy) and low-level decisions (e.g., where do I click next in the application to show cars with specific attributes?) while simultaneously representing relevant decisions that occur outside of those systems. We briefly review prominent and relevant typologies of the visualization community here and discuss why these typologies are insufficient for discussing decision-making.
Two of the most prominent and influential visualization task typologies in the community are the analytical task typology and the multi-level task typologies. The analytical tasks typology addressed a gap in highly data-centric typologies by taking an analytical-centric approach[3]. This typology consists of ten tasks that describe “primitive" analyses of data, or low-level tasks. The multi-level task typology further expanded upon these tasks to incorporate higher levels that can be composed from lower-level tasks, with the highest-level characterization of tasks falling into three categories of “why?", “how?", and “what?"[44, 13]. However, these tasks are not suitable for describing decision-making in compact yet meaningful ways. Consider the decision of buying a car by choosing among a set of available used cars. This task, in its current form, cannot yet be represented in either the analytical or multi-level task typologies because more information is needed on how the decision is made, which can vary across users. In other words, these typologies are oriented towards describing potential solutions or approaches for solving a problem, they do not describe the decision problem or goal. When we can describe decision problems in abstract yet precise ways, we can begin to reason about solutions more holistically.
3 Methodology
This section describes our methodology for developing a typology of decision-making tasks. Our approach began with gathering design requirements from a literature review of publications in the visualization community relating to decision-support tools. Specifically, we analyzed design studies and visual analytics application papers to extract example scenarios and relevant contexts relating to decision-making, as detailed in subsection 3.1.
In subsection 3.2, we explain how we formulated these design goals by examining the specific decisions that domain experts sought to achieve within their problem domains while using decision-support tools. This involved dissecting the decision processes, identifying key components, and highlighting other properties relevant to decision-making, as outlined in subsection 3.3.
Following this, we conducted collaborative working sessions with all authors to generate a comprehensive set of decision-making tasks that met the design goals. This was an iterative process that involved continuous refinement and validation of the decision tasks and their definitions, drawing upon prominent decision-making theory papers, as we discuss in subsection 3.4.
The result of this iterative design process is our proposed Typology of Decision-Making Tasks for Visualization, which we present in detail in Section 4.
3.1 Survey Corpus
We reviewed relevant visualization decision-support literature to formulate the design goals of our typology, building upon the literature corpora curated by Oral et al.[48] and Shi et al.[57]. Oral et al.’s corpus focused on papers explicitly labeled as decision-making tools, filtering based on the presence of the term ‘decision’ in titles or abstracts. However, this approach omitted valuable visualization papers which serve as decision-support tools for domain experts and tools implementing methodologies other than MCDM. To broaden the scope, we incorporated Shi et al.’s work, allowing us to encompass a more comprehensive range of visual analytics applications and design studies. From their list of 190 VAST papers[57], we found 63 to be applications with a direct decision-making focus. We added to that list the papers with “decision” mentioned in titles or abstracts to be consistent with[48], adding up to 74 papers.
To refine our selection, we prioritized papers featuring specific domain applications, and eliminated generic systems lacking evaluations with domain experts. This guaranteed that the chosen tools were utilized by domain users with actual decisions and requirements that could be validated. For instance, papers solely based on hypothetical case studies were excluded (e.g., Podium[63], Zooids[42], etc.). This filtering process resulted in a final corpus of 69 papers that directly contribute to our study’s objectives.
3.2 Coding Process
Our coding process followed an inductive approach to derive the codes, i.e. the design goals for our typology, from the survey corpus outlined in the preceding section. The process involved a team-based approach with two coders, who are the first two authors of this paper. For each paper, we identified the primary decision-making goal by examining the abstract and introduction. We identified a few potential goals and then determined the true goal by analyzing the task abstraction, requirements, evaluation sections, and other supplementary materials, such as figures, video demonstrations, and supplementary materials (if available). In addition, we generated codes for unique decision contexts, sub-decisions that supported the overall decision, and the information that those decisions would use and produce. The resulting codes are discussed in the subsequent subsection.
3.3 Design Goals for a Decision Task Typology
The generated codes were used to derive a set of design goals. Through multiple brainstorming sessions involving all authors, we distilled the following list of goals that we believe our typology should fulfill in order to accurately represent the decision processes described in the corpus’s case studies and decision-making scenarios. For each goal, we include excerpts from interviews that support our choice of that goal.
  • Iteration: Given that decision-making processes in visualization often involve cycles and iterations, our typology should incorporate a mechanism to express such iterative patterns.
    • By working with the domain experts, we constructed a threestage iterative workflow for analyzing and improving bus networks[71]
    • Planners, stakeholders, and development team members work together in planning their project using ReleasePlanner; the iterative standard planning procedures include…[7]
  • Hierarchical structure: Decisions are intricate and can be deconstructed into subtasks, necessitating the ability of our typology to capture these hierarchical relationships.
    • Our final design improved the layout and probability encoding of the hierarchical-oriented choices.[31]
    • We present two design lessons we have learned while developing BNVA. First, hierarchical exploration is mandatory for complex data like bus networks. […] In this study, we employed a hierarchical exploration approach to dissect a network with multiple coordinated views from the overview to the details and provide visual hints for the users to drill down in the dataset and find interesting patterns.[71]
  • Process-centric approach: Our typology should focus on delineating the sequence of decisions users encounter, rather than merely representing data-processing flow or interaction order typically outlined in task lists and design requirements of application papers and design studies.
    • Using a laboratory model, the authors have investigated various possible methods for breach closure, utilizing procedures such as single-and multi-barrier embankments with different ways of positioning sandbags. […] We have identified 7 major phases that led to the final solution.[68]
    • …capturing metadata of the visual exploration process is a key component of the analysis process. Our history tree visualization is able to record users’ decisions and allow them to compare, modify and insert new decisions.[1]
    • Our approach aims to implement the scenario pool concept as we see it: first, create many plans in the preparation phase; second, quickly pick the most appropriate one in the response phase.[70]
  • Data-type flexibility: The decision tasks within our typology should demonstrate flexibility and impartiality towards input and output types, instead emphasizing how information is processed and evaluated.
    • The person in charge [..] has an overview of an area where a disaster can happen, and a number of protection measures available. […] The goal of our steering environment is to allow the user to experiment with various options, and to do it as intuitively as possible.[53]
    • …only information on a few products is given to the buyers, leaving them wondering what other alternatives are available.[72]
    • This interactive exploratory tool helps the casual decision-maker quickly choose between various financial portfolio options and view possible outcomes.[54]
Additionally, we aim for our typology to possess the following desired properties:
  • Generalizability: The typology should apply uniformly across the diverse range of papers we surveyed and should remain relevant across various domains.
  • Simplicity: We aspire for our typology to remain straightforward and concise, with minimal decision types and operations to avoid unnecessary complexity.
  • Accessibility: Our typology should be easily understood and grasped by both the visualization designers and the domain experts, contributing to its usability and adoption.
These criteria guided the creation of our typology, as we discuss in the following subsection.
3.4 Iterative Design Process
Our iterative design process began with brainstorming an initial set of decision-making tasks, drawing from the design goals outlined in subsection 3.3. Each iteration involved consulting relevant decision-making theory literature, particularly works by Fischhoff et al.[28, 27], Simon’s decision making phases [58, 14], cognitive science and cognitive process literature[66, 51, 60, 22], to ensure alignment with established theories. In each iteration, we defined the newly introduced tasks and modified or deleted existing task definitions to eliminate overlaps.
Following this, we validated the set of decision tasks by testing them on the survey corpus, aiming to confirm their descriptiveness and properties highlighted in our list of design goals. During this validation phase, we noted any missing capabilities that required inclusion in subsequent iterations. Additionally, we identified tasks that were not utilized for removal in the next iteration.
The iterative process continued with modifications based on feedback from the validation step and revisiting relevant decision-making theory literature to ensure coherence and alignment. This process persisted until the set of decision-making tasks and their definitions reached stability, resulting in a finalized typology, which we describe below.
4 Typology of Decision-Making Tasks
As we outlined in the previous section, we created a typology that reflects and captures the ubiquity of decision-making while allowing for flexibility in the execution of decisions. Here we define the tasks of our typology, focusing on the unique properties of each. We then move to a discussion of the composability of the tasks, followed by four case studies that illustrate the outlined properties in real-world visualization systems.
4.1 Decision-Making Tasks
Our typology consists of three tasks derived from the scientific literature[28, 27]: CHOOSE, ACTIVATE, and CREATE. Each task is a function that represents a specific and distinct decision problem. The type of the inputs to these functions does not change the core process of the decision task. Some of the key differences between the tasks are the unique transformations of inputs into outputs, the existence of input features for evaluation, and the existence of well-defined evaluation functions.
4.1.1 CHOOSE
The CHOOSE task assesses options and outputs a subset deemed optimal or best. Specifically, CHOOSE takes as input a set of objects, and an evaluation function that determines a score for each input. Importantly, the evaluation of one input depends on the other inputs. In other words, whether an input is kept or selected for output depends on how well it scores relative to the other inputs.
Consider the example decision of buying a car; we have four cars from which we can CHOOSE. We could score the cars on their features relative to the other cars in the set, and select the car with the top score. Subsequent updates to the scoring function and the re-selection of the best car are represented as multiple executions of the CHOOSE task.
The core component of CHOOSE—the evaluation of inputs with respect to others—does not change, but there is flexibility in how evaluation is performed. As we discussed above, one way to execute the evaluation in a CHOOSE decision is to score each input, sort the alternatives according to that score, and select the top $k$. Alternatively, the evaluation could perform comparisons between each pair of inputs, keeping only the subset that wins in each comparison, and continue this process until $k$ inputs are left in the subset. In this case, no scoring function is required as long as there is a way to compare the inputs.
4.1.2 ACTIVATE
The ACTIVATE task represents a decision where inputs are evaluated, and only those that meet or exceed a threshold are returned. As in the previous task, ACTIVATE determines a score for each input using existing attributes on the inputs. However, the evaluation of one input is independent of the other inputs. Whether an input is returned in the output depends only on how well it scores, regardless of the scores of the other inputs.
In the car-buying example, the ACTIVATE task can be represented by applying thresholds to each existing feature and buying the car that satisfies all thresholds. For example, we could buy the car that is under a particular price and mileage and has a cargo capacity over a particular value. Notice that because each car is evaluated independently, we can assess them all in the evaluation simultaneously. Subsequent updates to the thresholds to eliminate all but a single car are represented as multiple executions of the ACTIVATE task.
Similar to the previous task, there is flexibility in the evaluation that occurs in ACTIVATE. For example, the threshold can change across inputs as a result of the features of each individual input. In the car buying decision, the price threshold may decrease if the mileage and cargo capacity are too close to the thresholds, or the price threshold could increase if the cargo capacity significantly exceeds the threshold. Importantly, this change in threshold does not depend on the other inputs or their features. Unlike CHOOSE, where evaluation does not require a scoring function because comparison or preference could be used instead, the ACTIVATE task must have one. This is because in ACTIVATE there is no other way to assess whether an input is included in the output.
4.1.3 CREATE
The CREATE task represents decisions on assembling, synthesizing, or generating new information. Specifically, CREATE can take any number of inputs (including none) and return outputs of a new type or augment inputs with new information. This is in contrast with the other tasks in our typology, where the inputs and outputs are consistent with each other. The evaluation in CREATE addresses how to assemble the inputs into outputs, or what outputs to produce given the inputs.
Consider again the decision to buy a car. Rather than deciding on an existing set of cars, the CREATE decision allows for three possibilities: (1) the creation of a new option that is not in the current set of options, (2) the addition (or removal) of a feature, and (3) synthesizing new information based on the inputs.
Of the three possibilities, the third use of CREATE is a departure from most decision-making literature and models, but is of particular importance to the visualization community. This decision represents the creation of knowledge or insights[9, 18, 38, 55], or more tangibly, reports[19], new data transformations[37], etc.
Similar to the other two tasks, the evaluation in CREATE can change according to the assessment of the input features. For example, we could consider building multiple possible cars, where each new car we add to the set is as different as possible from the others. In this case, the output of the CREATE task, the set of cars, is also an input alongside the set of features from which to build a new car to add to the set.
5 Composability
The decision tasks in our typology can be composed of or decomposed into other tasks in the typology. This is an important property that allows us to represent the hierarchies and complex structures that exist in real-world decision-making problems. For example, consider the decision to buy a car as a CHOOSE task. We could decompose this task into several sub-decision tasks like a CREATE decision which adds new possible cars to choose from, an ACTIVATE that filters cars that do not meet specific requirements, and a CHOOSE decision to select the best option. The ability to break down the tasks in our typology allows us to represent how smaller decisions help resolve the overall decision. This example illustrates two properties of our typology:
Composability: The three sub-decision tasks are interconnected to form a larger decision task. These sub-decisions are said to be composable because the output of one sub-decision task becomes the input of the next. For example, CREATE adds a decision critera, which is then used by ACTIVATE to filter out undesirable options. The property of composability is important for describing visualization tasks in general, as demonstrated by the task typology by Brehmer et al.[13]. The composition of these decision tasks can be visualized as a directed graph, referred to as a “diagram” throughout this paper. In these diagrams, directed edges represent the compositions of decision tasks. Importantly, when two tasks are connected by a directed edge, the input type of the task at the end of the edge must match the output type of the task at the beginning of the edge.
Hierarchy: The decomposition of the CHOOSE task in the above example into three sub-decision tasks implicitly defines a hierarchical structure of decision tasks. These hierarchies can be useful as they can describe a decision task at a high-level abstraction, while also offering detailed step-by-step decision-making processes at lower levels. While these hierarchies can be constructed either from the top-down or the bottom-up, during our interview study, we observed a preference among participants for a top-down approach (see Section 7). The participants started with high-level decision tasks and recursively decomposed them into sub-tasks until they achieved the desired level of granularity. In the case studies below, we demonstrate how these decision-task hierarchies can be created through a series of decompositions.
6 Case Studies
To illustrate the composability and hierarchical properties of our typology, we present four case studies where our typology is used to describe user decisions in the context of decision-support tools.
6.1 Case Study 1: Homefinder Revisited
Figure 2: Homefinder Revisited[72] decision diagram.
The paper “Homefinder Revisited”[72] introduces ReACH (Reachability-Aided Contemporary HomeFinder), a visual analytics system aimed at aiding users in discovering their ideal homes. In ReACH, users input their daily routes, activities, and preferences, upon which the system generates potential homes by considering factors like reachability, proximity to amenities, and user-specified preferences. With its interactive interface, users can query, filter, and evaluate home candidates based on multiple criteria, ultimately choosing their ideal home.
Level 1: At the root level, ReACH supports the user to CHOOSE their ideal home from a number of candidates.
Level 2: Expanding on the CHOOSE decision, a user needs to first CREATE home candidates from their routes, activities, and preferences. The user performs the ACTIVATE decision to add or filter the candidates by refining the search criteria (e.g., parameters for reachability, preferences, etc.). The user then decides the homes that they would like to add to their wishlist (an ACTIVATE decision) and iterates on the process until the wishlist is complete. Finally, the user CHOOSES the ideal home from the wishlist.
Discussion:
The decision-making framework described in the Homefinder Revisited paper represents a classic multi-criteria decision-making (MCDM) task. Typically, an MCDM task follows a diagram of CREATE (creating design criteria) $\rightarrow$ ACTIVATE (filtering and refining the criteria) $\rightarrow$ CHOOSE (selecting the best options). ReACH adds an additional decision task to this process, allowing users to add or remove candidates from a wishlist.
6.2 Case Study 2: Umbra
Figure 3: Umbra[65] decision diagram
Umbra[65] is a visual analytics tool that helps address the complex problem of sanitizing sensitive information within datasets without compromising data characteristics, particularly in scenarios involving medical data, personally identifiable information, and classified data. Umbra is semi-automated in that it can: (1) automatically identify potential risks, (2) suggest sanitization actions and schemes, and (3) perform simulated attacks. Using Umbra, an analyst would review the potential risks, run simulated attacks, adopt or develop a sanitization scheme, and finally execute the scheme on a given dataset and disseminate the result. The decision tasks supported in Umbra can be viewed hierarchically and are illustrated in Figure 3:
Level 1: At the root level, Umbra supports the ACTIVATE decision where a sanitized dataset is being evaluated on whether the changes made to it are acceptable or not, depending on its remaining utility and the privacy achieved.
Level 2: At the next level, the ACTIVATE decision can be expanded to include two additional decisions: a CHOOSE decision, where the users choose what sanitization scheme to apply to the dataset and where to apply it, and a CREATE decision, where a simulated attack is generated. This forms a loop: (1) choose a sanitization scheme, (2) run the simulation, and (3) accept or reject the sanitized data.
Level 3: The simple loop can be decomposed to include more nuanced decisions in the process. The CHOOSE decision of a sanitization scheme can be expanded into its own decision-flow that includes creating a sanitization approach (a CREATE decision), and deciding whether to accept the approach (an ACTIVATE decision). Similarly, the CREATE decision from Level 2 on generating a simulated attack can be decomposed further into deciding on a plan of attack (an ACTIVATE decision) and deciding whether to use custom or default test data (a CHOOSE decision).
Discussion:
In the first case study, the “Homefinder Revisited” paper describes the fundamental user task at the root level as a CHOOSE decision for choosing an ideal home from a finite list of candidates. However, in this second case study, the number of candidates is potentially infinite, as Umbra can generate as many sanitization schemes as required. As a result, the Umbra system is designed to assist users in examining and refining a scheme, culminating in the decision of whether the scheme can be ACTIVATED and applied to sanitize a dataset.
6.3 Case Study 3: Centaurus
Figure 4: Centaurus[25] decision diagram.
The Centaurus tool was utilized in a case study to evaluate the truthfulness of claims by analyzing related news articles[25]. Users were presented with a claim, such as a tweet, accompanied by a mix of trustworthy and untrustworthy news articles. Through Centaurus, users engaged with these articles to extract relevant information and assess the trustworthiness of each article. The primary focus was on understanding how users utilized Centaurus to formulate their final evaluations, irrespective of their stance on the original claim’s truthfulness. Figure 4 illustrates the following breakdown of the decision-making tasks:
Level 1: The core of the Centaurus decision problem is an ACTIVATE decision, wherein users evaluate the trustworthiness of articles.
Level 2: The Level 1 ACTIVATE decision is supported by additional decisions – an ACTIVATE to select relevant article based on search terms, a CHOOSE to choose articles for deeper analysis, and another ACTIVATE to search for similar articles after a user has made a decision about a potential article to investigate further.
Level 3: The ACTIVATE decision for selecting articles based on search terms is expanded to include a CREATE node where a user can manually enter keywords as search terms.
Discussion:
While the domains of Centaurus (determining article trustworthiness) and Umbra (choosing data sanitization) may seem different, they exhibit notable similarities in the decision-making processes. Figure 4 shows the decision diagram of Centaurus that uses a similar layout as the Umbra diagram in Figure 3. By comparing the two, one can observe that Centaurus features the same three tasks at the top row as Umbra (CREATE, ACTIVATE, CHOOSE), and the bottom row also shares similarities. This comparison highlights the utility of our typology in evaluating and contrasting the decision tasks of seemingly disparate decision-support visualization tools.
6.4 Case Study 4: WireVis
Figure 5: WireVis[16] decision diagram.
The WireVis paper introduces a visual analytics system designed for the analysis of large amounts of financial wire transactions, aiming to uncover suspicious or fraudulent activities such as money laundering[16, 17]. The use of the WireVis system entails two distinct yet interconnected decision tasks: (1) identify fraudulent transactions and their associated accounts, and (2) discover previously unknown strategies used to commit financial fraud. In both cases, analysts use a coordinated set of visualizations to examine financial transactions from multiple perspectives, including keywords, amounts, time, and account relations. The system can automatically generate hierarchical clusters and perform pattern matching, thereby assisting analysts in their investigations.
Focusing on the task of “discovering previously unknown strategies used to commit financial fraud,” the decision tasks can be represented as:
Level 1: The core objective can be described as a CREATE decision where an analyst aims to synthesize and discover a previously unknown “financial fraud strategy” based on suspicious financial transactions and activities.
Level 2: In order to CREATE the strategy, an analyst first need to CHOOSE the most promising accounts to investigate. For these accounts, the analyst determines the set of their transactions that might exhibit an unusual pattern of behavior (an ACTIVATE decision). These transactions are examined based on a list of criteria, such as the amount, the keyword used, the date patterns, etc. (an ACTIVATE decision). If deemed suspicious, an analyst might be able to deduce a previously unknown strategy of financial fraud (a CREATE decision).
Discussion:
The WireVis case study diverges from previous examples in that the root decision is not to simply choose or filter out potential options. Instead, it involves the creation of insights assembled from evidence gathered in the tool. The CREATE decision deviates from the typical multi-criteria decision-making (MCDM) style, and is a common and important aspect of decision-making within the visual analytics community.
7 Evaluation
Figure 6: Diagrams created by our participants during the interview study. Texts, labels, and annotations that contain identifiable information have been removed to preserve anonymity.
We recruited nine visualization experts to evaluate our typology from their perspectives and identify its strengths and weaknesses. Three of our participants are assistant professors, four are scientists at national laboratories, one is a senior visualization developer at a market research company, and one is a PhD student in visualization.
The interviews were semi-structured with a set of pre-determined open-ended questions with the opportunity to explore particular discussions further[36]. The interviews were conducted remotely via Zoom, each lasting approximately one hour, and recorded with the consent of all the participants. We asked the participants to choose a decision support tool they developed or an academic paper they wrote about during the interview. Then, we demonstrated our typology and its ability to compose decision-making tasks. This was followed by an exercise in which they had to describe the problem their tool addressed using our typology. After the exercise, we asked the participants open-ended questions about the properties of the typology. We concluded by asking them to reflect on other typologies and taxonomies in visualization and any suggestions for future work.
Overall, our interviews examined three broad areas: the typology’s correctness, utility, and comparison with existing task taxonomies and typologies commonly used in the visualization community. We address each of these areas in the following subsections.
7.1 Correctness
The participants discussed two topics related to the correctness of the typology: (1) how complete is the typology, and (2) how well the participants were able to use the typology to express their decision-making problems.
7.1.1 Completeness
In evaluating the correctness of our typology of decision-making tasks for visualization, we first assessed its completeness to determine if there were any missing decision tasks in our typology that could limit its ability to capture all aspects of decision-making. The participants in our study found that our typology was indeed comprehensive, allowing them to represent all decisions within the problem domain. As noted by one of the participants: “I think [the typology] is a pretty good set of high level operations. […] It took me a little bit to really clarify the distinction between activate and choose. But once I got there, it makes a lot of sense. I don’t know if I can really think of another type of operation that I felt was missing. It seemed, at least in the case of my problem, a good set [of operations] to describe what’s going on.” Other participants expressed similar sentiments. Overall, all of the participants were able to describe their decision-making problems using our typology, and none of them identified any missing decision tasks or fundamental operations in our typology.
Figure 7: A hierarchy annotation used by one of the participants during the interview study. In this figure, the triangles (and the arrows) indicate the role of a decision-maker in an organization’s hierarchy.
Interestingly, a few participants incorporated additional annotations when describing their decision-support tools using our typology (e.g., see Figure 7). These annotations were used to detail the contexts of these participants’ decision-making problems, including temporal information (e.g., time duration for how long a task would take or how frequently a task needs to be performed), roles (e.g., who in an organizational hierarchy should make the decision. Or, in a collaborative setting, which collaborator is responsible for which task), and decision styles (e.g., rational, conceptual, analytical, etc.). Such contextual annotations complement our typology and suggest an interesting future extension of our work. We will explore these concepts in greater detail in Section 8.
7.1.2 Expressiveness
Participants unanimously expressed their proficiency in composing and decomposing decisions with the typology, enabling them to construct cohesive representations of the problems their respective visualization tools addressed. They noted that the typology adeptly captured the decision-making processes inherent in problem domains and facilitated the visualization of user interactions with decision-support tools. As one of the participants said, “[The use of the typology] really kind of helped highlight some of the complexity around the design goals that went into it. And it’s definitely reflective of conversations that [my team and I] had going into the design of the tool and sort of throughout its life span. So, in that sense, I think it’s very representative.
This sentiment is perhaps best summarized by a participant: “Breaking down this higher level activate task…in terms of how [the user] actually interacts with the system and what the system is doing for them, I think [the typology] helped tremendously. […] I was able to describe step by step what people were doing in the system, and those steps mapped really cleanly to the [three decision-making] tasks.
7.2 Utility in the design process
During the interviews, the participants discussed the potential applications of the typology. We categorize their feedback into three themes: the use of the topology for (1) analysis and reflection for guiding the design process of visualization tools, (2) facilitating communication between developers, domain experts, and end-users, and (3) categorizing and organizing VIS publications related to decision-making.
7.2.1 Analysis and Reflection of Design
Participants reported that structuring decisions in a graph or hierarchical format using our typology helped them break down problems and decision goals, identify potential decisions that were not initially considered, and compare the current designs with improvements. In particular, the graph served as an externalization of mental models for the participants to better reflect on the overall decision-making process, allowing them to step back from the development of tools and consider the bigger picture. As explained by one of the participants, “I think this [exercise using the typology] was actually very valuable from my standpoint in reflecting on what we are building. Take a step back and reflecting on how these decisions, what we understand, and from an abstraction and typology standpoint, what they mean. So it’s been actually helpful in that regard.”
Another participant further delineated the benefits of using the typology, “I see this as a way […] for us to understand clearly what we are doing. Not to focus a lot on how we do it, but what we want to do. […] Breaking down the task like that could help me to also think about new possibilities for decision-making that I didn’t consider before or didn’t have time to.
7.2.2 Facilitate Communication
During our interviews, participants highlighted the typology’s potential to facilitate communication with domain experts and end-users during the design process by providing a higher-level perspective. One participant articulated, “If I had this taxonomy like in your work, I can better communicate with the experts. […] I think with this taxonomy maybe we can begin from a more high level perspective, like what does the system look like or how this system would adjust its requirements from a higher perspective. And then we can drill down to each part, for creation, for activation, for choose. What kind of interface we can design for them. I think [this diagram] would facilitate this kind of brainstorming sessions.
By facilitating better communication, a participant believed that the use of the typology could improve shared understanding of the design process between the stakeholders, thereby improving the final design. “I think having some kind of decision-making type of language might help bridge that gap between ‘I as the developer, designer, researcher, I think I know what you mean, I think I know what you want, and what your processes are’ as opposed to the user being like ‘I don’t even know how to better explain this.”’ Similarly, another participant compared the use of the typology to the standard practice that relies on a list of design requirements in the design process and noted, “I think this [diagram] is a better specification of the process than the list of requirements, for sure, because at least I can see here the process of the decision, not a well laid set of requirements.
Although our evaluation did not involve domain experts or end-users, our participants speculated that our typology would be intuitive for their domain experts and end-users to understand, thereby improving the design process. One participant noted, "I don’t think you could really present [Brehmer et al.’s task typology] to someone outside of the [visualization] community and explain it in 10 minutes and they could like really, intuitively, get it. [Your typology] feels much more accessible." Another participant said, “I haven’t practically used the diagram [with a domain expert, but] I think the expert could provide feedback directly on the diagram. Basically, if they want more alternatives, or if they think that the current way of generating alternatives is not satisfactory, they can request to add edge backward to the creation tasks. So our system design will improve.” This direct engagement with the diagram has the potential to improve the collaborative refinement of system design, allowing experts to contribute insights that could enhance decision-making capabilities.
7.2.3 Categorization and Organization of Publications
Some participants highlighted the potential of our typology in aiding the research community in categorizing and organizing publications relating to decision-making effectively. In our interviews, three participants utilized our typology to sketch diagrams representing different versions of their tools or extensions published in subsequent papers, demonstrating its capacity to compare decision processes across various tools. Through these diagrams, they were able to discuss the nuances and distinctions between these tools, highlighting areas with varying levels of decision support. This process enabled them to evaluate and compare their own tools’ effectiveness.
The participants expressed that our typology could serve as a valuable tool for evaluating decision support tools due to its ability to organize the decision process into distinct parts and delineate where decisions begin and end. One participant noted, “In my experience as a reviewer, I don’t explicitly ask authors to follow specific patterns or taxonomies to organize their work. However, if they use a taxonomy, it will be easier to follow their paper. Basically, they can organize decision-making papers into these few stages, making it easier to evaluate this type of work.”
Instead of delving into the minutiae of individual visualizations in the review process, one participant suggested that our typology could offer a framework to assess systems holistically, stating that, “I think [your typology] could be helpful as well when we evaluate visualizations. [In reviewing visualizations] we go very low level [evaluating if] people [can] read a certain value [from a visualization]. […] [T]hinking about [a decision support] system [using your typology], like in this way, might be also helpful to evaluate the systems overall, because [as a reviewer] I know where I start and where I want to end. And I can see how [a user would] follow these steps to get to the final process.
7.3 Comparison with Existing Task Taxonomies
All of our participants are researchers in the VIS community who are well-versed in the VIS literature. Many drew comparisons between our typology and existing task taxonomies, such as those proposed by Amar et al.[3] and Brehmer et al.[13]. In this section, we summarize their comments on how our typology compares to existing task taxonomies in terms of: (1) the level of task abstraction and (2) its flexibility in practical use.
7.3.1 Level of Task Abstraction
During discussions, participants noted that our typology is more oriented towards modeling processes in contrast to the focus on data, analytical operations, and interactions emphasized by existing task taxonomies. This approach resonated with them as they considered it more suitable for describing complex decision-making tasks. One participant articulated this sentiment as follows: “I think [existing task taxonomies] can be suitable for all visual analytics systems and it’s a bit low level […]. [But] if you want to describe my system with this type of taxonomies you have to go very low level, like presenting values or like sorting the alternatives. […] It’s a bit cumbersome to describe the system using this taxonomy. Yeah, yours are better. I think your taxonomy is doing a better job at describing decision-making tasks or decision-making systems, because it’s from the number of alternative perspective. So I think that the taxonomy more fits in a decision-making scenario.
Other participants also noted the challenge of using existing task taxonomies to describe a complex design-support visualization system and suggested that a higher-level typology is better suited for describing the user’s needs and perspectives. As stated by one of the participants, “[Using existing taxonomies] is very low level to understand a system like [my system]. For the final users, the user is not thinking about reading this bar chart. [The] users say ‘can I choose my actual scenario?”’ Another participant agreed, "[Your typology] is abstract, but it helps you think about the decision process. This is closer to the user than the why of a visualization [in the typology by Brehmer et al.], because you don’t specify, for example, ’I want to present data,’ or ’I want to discover data.’ […] [With a] diagram [using your typology], […] I can see a process that I can match well with what we did in [our system].
7.3.2 Flexibility to Data and Context
The participants in our evaluation highlighted the flexibility of our typology, particularly in its adaptability to different data and contexts while maintaining well-specified functions. The participants could compose and decompose decisions according to their decision tasks and preferences for modeling the problems. For example, one participant chose to represent their tool’s addressed problem hierarchically, noting, “This does feel more hierarchical. I’m definitely reminded of task abstraction type of work, which tends to go a step beyond just requirements gathering.
Other participants opted for the diagram representation, utilizing cycles and arrows to depict iterations within their decision-making process. This approach allowed for clear visualization of feedback loops, with decisions feeding back into the initial stages of the process. One participant who adopted this method remarked, "[Brehmer et al.’s typology] very much still had kind of a rigid application to it, whereas I think the typology that you’re proposing feels like it can be more flexible."
The general agreement among participants was that our typology stands out for its flexibility, allowing for diverse representations of decision-making processes tailored to specific needs and contexts. A participant summarized this quite succinctly by stating, “I like that a lot as part of your typology, is that it feels simple and flexible, yet powerful.
8 Discussion & Future Work
The evaluation of our decision-making task typology received predominantly positive feedback from the participants in our interview study. In this section, we analyze additional feedback from participants to identify potential areas for future research.
8.1 Annotations in the Typology
During the interview study, several participants incorporated annotations into their diagrams or requested such annotations while illustrating their decision-support tools using our typology. These annotations did not expand the set of tasks in our typology or modify their definitions, and they provided additional context for the decisions.
The annotations our participants added included the frequency of decision-making instances, the roles of individuals in a collaborative or organizational setting, and decision-making styles. For instance, one participant highlighted the potential use of our typology to depict decision-making problems based on various “personas,” resulting in distinct diagrams reflecting the roles of decision-makers within an organization. Another participant proposed annotating the diagram to indicate which decision tasks humans perform and which involve computer assistance or collaboration.
Ultimately, the incorporation of annotations highlights an opportunity to further examine decision-making in visualization with our typology. For instance, “personas" could be described through different diagrams for solving the same decision problem. Alternatively, denoting decision tasks performed by humans or computers allows us to concretely express and investigate the role of artificial intelligence in decision-making. Our typology can facilitate deeper exploration and reflection on the many facets that intersect with decision-support visualization systems.
8.2 Measuring the Effectiveness of the Typology
In addition to conducting an interview study with visualization researchers to evaluate our typology, we also assess its effectiveness according to the criteria established by Brehemer et al.[13] and Yi et al.[74]. Specifically, we examine our typology through the lens of three dimensions proposed by Beaudouin-Lafon[10]: (1) descriptive power: “the ability to describe a significant range of existing [designs]”, (2) evaluative power: “the ability to help assess multiple design alternatives”, and (3) generative power: “the ability to help designers create new designs.”
We have showcased the descriptive capabilities of our typology through four case studies and an interview study. In each instance, designers successfully utilized our typology to construct diagrams that effectively described their decision-making problems. Furthermore, we have begun to demonstrate how the typology could facilitate the evaluation and the identification of similar decision problems. Our case studies illustrated the potential of using these diagrams to compare the similarities and differences between domain specific decision problems in a domain-agnostic way.
Finally, we assess the generative aspect of our typology using the definition provided by Brehemer et al.[13]. Specifically, a typology is generative if has a systematic approach for using it to create workflows or diagrams. Our proposed hierarchical decomposition process of decision tasks corresponds with this definition. Starting from the root level, participants demonstrated the ability to recursively expand each node into a composed set of decision tasks, culminating in the creation of a complete diagram. This approach allows visualization designers to methodically externalize decision problems in a structured manner, facilitating a gradual but precise representation of the decision-making process.
8.3 From Typology to Visualization Design
Some participants in our study highlighted the opportunity to integrate the diagrams created using our typology with visualization designs. By pairing these diagrams with their associated visualization designs, it becomes feasible to explore these relationships to develop recommendation systems of decision-support tools, following a similar approach as the Draco system[43]. For instance, the Homefinder Revisited case study, as described in Section 6.1, illustrates a simple diagram representing a multi-criteria decision-making (MCDM) problem. Due to the widespread use of MCDM, this motif occurs in many other, more intricate decision-making scenarios. In the Umbra and Centaurus case studies illustrated in Figures 3 and 4, the same MCDM motif is evident (observed in the top three nodes), indicating that MCDM is a component of both the Umbra and Centaurus decision problems as well. When paired with their associated visualization designs, we can begin to identify and extract visualization design patterns that represent MCDM or other similar sub-decision tasks. In our future work, we aim to investigate how we can integrate our typology diagrams with visualization design processes to assist designers of decision-support tools.
9 Conclusion
This paper presents a typology for decision-making tasks in visualization, addressing the limitations of existing taxonomies. Built upon prior research and informed by design goals derived from a thorough literature review, the typology comprises three decision tasks: CHOOSE, ACTIVATE, and CREATE. These tasks allow for the representation of complex decision-making structures, as they can be composed or decomposed into other tasks. The typology demonstrates completeness, expressiveness, and utility across various decision-making contexts and is validated through semi-structured interviews with visualization experts.
In summary, this paper contributes potential insights into decision-making in visualization, offering a typology that can be used for analyzing design, facilitating communication, and categorizing visualization publications on decision-making systems. By providing a structured approach to understanding decision-making tasks, the typology offers a tool for researchers and practitioners alike, with implications for improving designs of decision-support tools in the future.